# Unifying Unsupervised Graph-Level Anomaly Detection and Out-of-Distribution Detection: A Benchmark

### v1

@@ Appendix B Detailed Description of Algorithms in UB-GOLD
Weisfeiler-Leman Subtree Kernel (WL) [31] : This kernel generates graph embeddings by iteratively refining node labels based on subtree patterns. It effectively captures structural similarities within graphs, making it a powerful tool for embedding generation.
Propagation Kernel (PK) [21]: This method propagates labels through the graph structure, resulting in embeddings that reflect the graph’s topology. It captures relational information within the graph, providing a robust basis for subsequent outlier detection.
@@
GraphCL (GCL) [32]: GCL leverages augmentations to learn robust graph-level representations. By contrasting different views of the same graph, the model learns to capture essential graph structures and properties, making it highly effective for various graph-based tasks. This method is known for its strong performance in unsupervised learning settings.
InfoGraph (IG) [33]: IG maximizes the mutual information between local and global graph representations to achieve unsupervised and semi-supervised graph-level representation learning. By capturing meaningful information across different levels of the graph, IG effectively learns comprehensive graph embeddings that are useful for a wide range of downstream tasks.
@@
Isolation Forest (iF) [23]: This algorithm isolates observations by constructing an ensemble of trees, each built by randomly selecting features and split values. An anomaly score is calculated based on the path length from the root to the leaf node, with shorter paths indicating anomalies.
One-Class SVM (OCSVM) [22]: OCSVM aims to separate the data from the origin using a hyperplane in a high-dimensional space. Points that lie far from the hyperplane are considered outliers.
@@
OCGIN [24]: Utilizes a GIN encoder optimized with a Support Vector Data Description (SVDD) objective to identify anomalies within the graph structure. This method employs an end-to-end GNN model with one-class classification for effective anomaly detection.
GLocalKD [3]: Jointly learns two GNNs and performs graph-level and node-level random knowledge distillation between their learned representations. By leveraging both local and global knowledge, this approach enhances the detection of anomalies.
OCGTL [7]: Extends deep one-class classification to a self-supervised detection approach using neural transformations and graph transformation learning as regularization. This technique improves the model’s unsupervised anomaly detection capabilities.
SIGNET [8]: Proposes a self-interpretable graph-level anomaly detection framework that infers anomaly scores while providing subgraph explanations. By maximizing mutual information of multi-view subgraphs, it achieves both detection and interpretation of anomalies.
GLADC [27]: Incorporates graph-level adversarial contrastive learning to identify anomalies. Through the creation of adversarial examples and learning robust representations, this method effectively distinguishes between normal and abnormal graphs.
CVTGAD [28]: Introduces a simplified transformer with cross-view attention for unsupervised graph-level anomaly detection. It overcomes the limited receptive field of GNNs by using a transformer-based module to capture relationships between nodes and graphs from both intra-graph and inter-graph perspectives.
@@
GOOD-D [9]: Performs perturbation-free graph data augmentation and utilizes hierarchical contrastive learning on the generated graphs for graph-level OOD detection. By leveraging multiple levels of contrastive learning, GOOD-D enhances the representation learning process, making it robust in distinguishing between in-distribution (ID) and out-of-distribution (OOD) graphs.
GraphDE [10]: Models the generative process of the graph to characterize distribution shifts. Using variational inference, GraphDE infers the environment from which a graph sample is drawn. This generative approach effectively identifies ID and OOD graphs by detecting shifts in the underlying data distribution.
Appendix C Supplemental Information of UB-GOLD
C.1 Metrics
@@
C.2 Additional Experimental Details
Implementation Details. To ensure a comprehensive evaluation and maintain fairness across a broad spectrum of models, we develop an open-source toolkit named UB-GOLD. This toolkit is built on top of Pytorch 2.01 [49], torch_geometric 2.4.0 [50] and DGL 2.1.0 [51]. We implement graph kernel methods with the DGL library. All other models are unified using the torch_geometric library. GCL and IG are included via the PYGCL library [52].
@@ C.2 Additional Experimental Details
C.3 New Experiments Dataset Split
In the experiments for generalizability and robustness, we do not follow the original split but utilize experiment-specific splits for fair comparison. The details are as follows.
Generalizability on Near-OOD and Far-OOD (Sec. 4.2). To rigorously evaluate the performance of GLOD and GLAD methods in handling near and far OOD conditions, we have implemented a distinct partitioning strategy for this research question. Unlike the setup in RQ1, here we ensure that the number of samples in the Near OOD, Far OOD, and ID Test groups are precisely equal, detailed in Table 5. This balanced configuration is designed to provide a fair comparison across different degrees of OOD scenarios, and it includes two specific setups:
@@ Appendix D Reproducibility
Datasets. Our datasets are publicly available and include TUDataset, OGB, TOX21, DrugOOD, and GOOD. Among them, TUDataset [30], OGB [53], and TOX21 [29] are licensed under the MIT License. DrugOOD [12] is licensed under the GNU General Public License 3.0. GOOD [19] is licensed under GPL-3.0.
@@
Near OOD | 50.09 | 71.87 | 100.00 | 49.41 | 50.61 | 97.65 | 40.30 | 47.18 | 100.00 | 56.61 | 57.25 | 98.00 |
Appendix E Additional Experimental Results
@@
Main experiments. Table 8 presents the complete results of the main experiment. Each model is executed 5 times with varying random seeds, and the mean scores along with standard deviations are reported. “Avg. AUROC”, “Avg. FPR95”, “Avg. AUPRC”, and “Avg. Rank” indicate the average AUROC, FPR95, AUPRC, and rank across all datasets, respectively. We observe that the End-to-End approach stands out in AUROC, achieving the best overall results. However, in the metrics of AUPRC and FPR95, SSL-D in the 2-step approach also performs well, showing competitive results.
@@
GK-D (2-step) | SSL-D (2-step) | GNN-based GLAD (E2E) | GNN-based GLOD (E2E) | |||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | |
p53 | ||||||||||||||||
HSE | ||||||||||||||||
MMP | ||||||||||||||||
PPAR | ||||||||||||||||
COLLAB | ||||||||||||||||
IMDB-B | ||||||||||||||||
REDDIT-B | ||||||||||||||||
ENZYMES | ||||||||||||||||
PROTEINS | ||||||||||||||||
DD | ||||||||||||||||
BZR | ||||||||||||||||
AIDS | ||||||||||||||||
COX2 | ||||||||||||||||
NCI1 | ||||||||||||||||
DHFR | ||||||||||||||||
IM→IB | ||||||||||||||||
EN→PR | ||||||||||||||||
AI→DH | ||||||||||||||||
BZ→CO | ||||||||||||||||
ES→MU | ||||||||||||||||
TO→SI | ||||||||||||||||
BB→BA | ||||||||||||||||
PT→MU | ||||||||||||||||
FS→TC | ||||||||||||||||
CL→LI | ||||||||||||||||
HIV-Size | ||||||||||||||||
ZINC-Size | ||||||||||||||||
HIV-Scaffold | ||||||||||||||||
ZINC-Scaffold | ||||||||||||||||
IC50-Size | ||||||||||||||||
EC50-Size | ||||||||||||||||
IC50-Scaffold | ||||||||||||||||
EC50-Scaffold | ||||||||||||||||
IC50-Assay | ||||||||||||||||
EC50-Assay | ||||||||||||||||
Avg. AUROC | 52.69 | 53.67 | 57.56 | 52.91 | 52.11 | 50.35 | 61.29 | 61.50 | 66.00 | 64.29 | 73.14 | 75.86 | 65.50 | 71.07 | 71.05 | 60.38 |
Avg. Rank | 12.09 | 11.94 | 10.46 | 11.77 | 12.06 | 11.65 | 8.14 | 8.31 | 6.69 | 8.23 | 4.91 | 3.34 | 7.14 | 4.51 | 4.54 | 10.14 |
GK-D (2-step) | SSL-D (2-step) | GNN-based GLAD (E2E) | GNN-based GLOD (E2E) | |||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | |
p53 | ||||||||||||||||
HSE | ||||||||||||||||
MMP | ||||||||||||||||
PPAR | ||||||||||||||||
COLLAB | ||||||||||||||||
IMDB-B | ||||||||||||||||
REDDIT-B | ||||||||||||||||
ENZYMES | ||||||||||||||||
PROTEINS | ||||||||||||||||
DD | ||||||||||||||||
BZR | ||||||||||||||||
AIDS | ||||||||||||||||
COX2 | ||||||||||||||||
NCI1 | ||||||||||||||||
DHFR | ||||||||||||||||
IM→IB | ||||||||||||||||
EN→MU | ||||||||||||||||
AI→DH | ||||||||||||||||
BZ→CO | ||||||||||||||||
ES→MU | ||||||||||||||||
TO→SI | ||||||||||||||||
BB→BA | ||||||||||||||||
PT→MU | ||||||||||||||||
FS→TC | ||||||||||||||||
CL→LI | ||||||||||||||||
HIV-Size | ||||||||||||||||
ZINC-Size | ||||||||||||||||
HIV-Scaffold | ||||||||||||||||
ZINC-Scaffold | ||||||||||||||||
IC50-Size | ||||||||||||||||
EC50-Size | ||||||||||||||||
IC50-Scaffold | ||||||||||||||||
EC50-Scaffold | ||||||||||||||||
IC50-Assay | ||||||||||||||||
EC50-Assay | ||||||||||||||||
Avg. AUPRC | 43.27 | 40.20 | 46.52 | 45.51 | 47.85 | 48.53 | 55.11 | 54.87 | 62.15 | 56.96 | 64.98 | 66.24 | 56.21 | 61.77 | 62.54 | 49.97 |
Avg. Rank | 11.26 | 12.26 | 11.34 | 11.06 | 10.89 | 10.03 | 7.86 | 8.14 | 6.20 | 8.66 | 5.63 | 3.77 | 7.77 | 5.31 | 5.34 | 10.49 |
GK-D (2-step) | SSL-D (2-step) | GNN-based GLAD (E2E) | GNN-based GLOD (E2E) | |||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | |
p53 | ||||||||||||||||
HSE | ||||||||||||||||
MMP | ||||||||||||||||
PPAR | ||||||||||||||||
COLLAB | ||||||||||||||||
IMDB-B | ||||||||||||||||
REDDIT-B | ||||||||||||||||
ENZYMES | ||||||||||||||||
PROTEINS | ||||||||||||||||
DD | ||||||||||||||||
BZR | ||||||||||||||||
AIDS | ||||||||||||||||
COX2 | ||||||||||||||||
NCI1 | ||||||||||||||||
DHFR | ||||||||||||||||
IM→IB | ||||||||||||||||
EN→PR | ||||||||||||||||
AI→DH | ||||||||||||||||
BZ→CO | ||||||||||||||||
ES→MU | ||||||||||||||||
TO→SI | ||||||||||||||||
BB→BA | ||||||||||||||||
PT→MU | ||||||||||||||||
FS→TC | ||||||||||||||||
CL→LI | ||||||||||||||||
HIV-Size | ||||||||||||||||
ZINC-Size | ||||||||||||||||
HIV-Scaffold | ||||||||||||||||
ZINC-Scaffold | ||||||||||||||||
IC50-Size | ||||||||||||||||
EC50-Size | ||||||||||||||||
IC50-Scaffold | ||||||||||||||||
EC50-Scaffold | ||||||||||||||||
IC50-Assay | ||||||||||||||||
EC50-Assay | ||||||||||||||||
Avg. FPR95 | 94.02 | 92.76 | 88.88 | 93.54 | 89.39 | 89.02 | 77.13 | 76.75 | 75.66 | 76.60 | 66.40 | 69.62 | 74.59 | 73.72 | 75.02 | 91.23 |
Ave. Rank | 11.60 | 10.54 | 11.63 | 11.23 | 10.23 | 9.94 | 7.60 | 6.54 | 7.46 | 8.74 | 6.04 | 4.21 | 6.56 | 6.28 | 6.41 | 10.74 |

### v2

@@ Appendix C Detailed Description of Algorithms in UB-GOLD
Weisfeiler-Leman Subtree Kernel (WL) (Li et al., 2016) : This kernel generates graph embeddings by iteratively refining node labels based on subtree patterns. It effectively captures structural similarities within graphs, making it a powerful tool for embedding generation.
Propagation Kernel (PK) (Neumann et al., 2016): This method propagates labels through the graph structure, resulting in embeddings that reflect the graph’s topology. It captures relational information within the graph, providing a robust basis for subsequent outlier detection.
@@
GraphCL (GCL) (You et al., 2020): GCL leverages augmentations to learn robust graph-level representations. By contrasting different views of the same graph, the model learns to capture essential graph structures and properties, making it highly effective for various graph-based tasks. This method is known for its strong performance in unsupervised learning settings.
InfoGraph (IG) (Sun et al., 2020): IG maximizes the mutual information between local and global graph representations to achieve unsupervised and semi-supervised graph-level representation learning. By capturing meaningful information across different levels of the graph, IG effectively learns comprehensive graph embeddings that are useful for a wide range of downstream tasks.
@@
Isolation Forest (iF) (Liu et al., 2008): This algorithm isolates observations by constructing an ensemble of trees, each built by randomly selecting features and split values. An anomaly score is calculated based on the path length from the root to the leaf node, with shorter paths indicating anomalies.
One-Class SVM (OCSVM) (Amer et al., 2013): OCSVM aims to separate the data from the origin using a hyperplane in a high-dimensional space. Points that lie far from the hyperplane are considered outliers.
@@
OCGIN (Zhao & Akoglu, 2023): Utilizes a GIN encoder optimized with a Support Vector Data Description (SVDD) objective to identify anomalies within the graph structure. This method employs an end-to-end GNN model with one-class classification for effective anomaly detection.
GLocalKD (Ma et al., 2022): Jointly learns two GNNs and performs graph-level and node-level random knowledge distillation between their learned representations. By leveraging both local and global knowledge, this approach enhances the detection of anomalies.
OCGTL (Qiu et al., 2022): Extends deep one-class classification to a self-supervised detection approach using neural transformations and graph transformation learning as regularization. This technique improves the model’s unsupervised anomaly detection capabilities.
SIGNET (Liu et al., 2023b): Proposes a self-interpretable graph-level anomaly detection framework that infers anomaly scores while providing subgraph explanations. By maximizing mutual information of multi-view subgraphs, it achieves both detection and interpretation of anomalies.
GLADC (Luo et al., 2022): Incorporates graph-level adversarial contrastive learning to identify anomalies. Through the creation of adversarial examples and learning robust representations, this method effectively distinguishes between normal and abnormal graphs.
CVTGAD (Li et al., 2023): Introduces a simplified transformer with cross-view attention for unsupervised graph-level anomaly detection. It overcomes the limited receptive field of GNNs by using a transformer-based module to capture relationships between nodes and graphs from both intra-graph and inter-graph perspectives.
@@
GOOD-D (Liu et al., 2023a): Performs perturbation-free graph data augmentation and utilizes hierarchical contrastive learning on the generated graphs for graph-level OOD detection. By leveraging multiple levels of contrastive learning, GOOD-D enhances the representation learning process, making it robust in distinguishing between in-distribution (ID) and out-of-distribution (OOD) graphs.
GraphDE (Li et al., 2022): Models the generative process of the graph to characterize distribution shifts. Using variational inference, GraphDE infers the environment from which a graph sample is drawn. This generative approach effectively identifies ID and OOD graphs by detecting shifts in the underlying data distribution.
AAGOD (Guo et al., 2023): A data-centric framework for graph OOD detection that operates on well-trained GNNs without retraining. AAGOD uses an Adaptive Amplifier to modify input graphs, highlighting key patterns for OOD detection. Through its Learnable Amplifier Generator (LAG) and Regularized Learning Strategy (RLS), it significantly improves detection performance and efficiency.
GOODAT (Wang et al., 2024a): A test-time graph OOD detection method designed to operate on well-trained GNNs without requiring training data or altering the GNN architecture. GOODAT employs a graph masker and the Graph Information Bottleneck (GIB) principle to extract informative subgraphs. Utilizing GIB-boosted loss functions effectively distinguishes between ID and OOD graphs, achieving strong performance across diverse datasets.
Appendix D Supplemental Information of UB-GOLD
D.1 Definition of GLAD and GLOD
GLAD definition. For GLAD, the task is to detect anomalous graphs within a given distribution . The objective is to assign anomaly scores such that anomalous graphs are distinguishable from normal graphs based on a threshold :
The theoretical objective for the anomaly scoring function is to maximize the separation between the distributions of normal and anomalous graphs:
where and represent the idealized distributions of normal and anomalous graphs. While this objective cannot be directly optimized during training due to the lack of access to test distributions, it provides a guiding principle for designing . Practical approaches typically rely on approximations or surrogate objectives based on training data, such as unsupervised or self-supervised learning frameworks, to approximate the separation of score distributions. When the separation is achieved, a threshold can effectively classify graphs as normal () or anomalous (), aligning with the GLOD framework.
GLOD definition. For GLOD, the task is to detect graphs in a test dataset that originate from a different distribution compared to the training distribution . The goal is to assign OOD scores such that OOD graphs have significantly lower scores than ID graphs, enabling a threshold-based classification.
The theoretical objective for the OOD scoring function is to maximize the separation between the distributions of ID and OOD graphs:
where and represent the idealized distributions of ID and OOD graphs, respectively. Although these distributions are not directly accessible during training (since OOD samples are not present in the training phase), they provide a theoretical goal for designing . In practice, approximations are used based on the available training data to model the OOD detection function. When the distributions of ID and OOD scores are well-separated, a threshold can effectively classify graphs as ID () or OOD ().
D.2 Metrics
@@
D.3 Additional Experimental Details
Implementation Details. To ensure a comprehensive evaluation and maintain fairness across a broad spectrum of models, we develop an open-source toolkit named UB-GOLD. This toolkit is built on top of Pytorch 2.01 (Paszke et al., 2019), torch_geometric 2.4.0 (Fey & Lenssen, 2019) and DGL 2.1.0 (Wang et al., 2019). We implement graph kernel methods with the DGL library. All other models are unified using the torch_geometric library. GCL and IG are included via the PYGCL library (Zhu et al., 2021).
@@ D.3 Additional Experimental Details
D.4 New Experiments Dataset Split
In the experiments for OOD sensitivity spectrum and robustness, we do not follow the original split but utilize experiment-specific splits for fair comparison. The details are as follows.
OOD sensitivity spectrum on Near-OOD and Far-OOD (Sec. 4.2). To rigorously evaluate the performance of GLOD and GLAD methods in handling near and far OOD conditions, we have implemented a distinct partitioning strategy for this research question. Unlike the setup in RQ1, here we ensure that the number of samples in the Near OOD, Far OOD, and ID Test groups are precisely equal, detailed in Table 5. This balanced configuration is designed to provide a fair comparison across different degrees of OOD scenarios, and it includes two specific setups:
@@ Appendix E Reproducibility
Datasets. Our datasets are publicly available and include TUDataset, OGB, TOX21, DrugOOD, and GOOD. Among them, TUDataset (Morris et al., 2020), OGB (Hu et al., 2020), and TOX21 (Abdelaziz et al., 2016) are licensed under the MIT License. DrugOOD (Ji et al., 2023) is licensed under the GNU General Public License 3.0. GOOD (Gui et al., 2022) is licensed under GPL-3.0.
@@
Near OOD | 50.09 | 71.87 | 100.00 | 49.41 | 50.61 | 97.65 | 40.30 | 47.18 | 100.00 | 56.61 | 57.25 | 98.00 | |
AAGOD | Far OOD | 94.95 | 94.80 | 8.25 | 78.39 | 75.42 | 21.50 | 57.92 | 64.35 | 87.00 | 54.23 | 60.45 | 94.00 |
Near OOD | 84.32 | 83.45 | 18.75 | 60.08 | 57.80 | 42.34 | 56.34 | 58.72 | 88.00 | 51.60 | 55.23 | 96.00 | |
GOODAT | Far OOD | 93.28 | 92.88 | 10.00 | 82.43 | 80.32 | 24.11 | 65.97 | 71.25 | 89.00 | 53.78 | 60.25 | 94.50 |
Near OOD | 81.93 | 80.75 | 22.50 | 62.92 | 61.23 | 45.36 | 56.98 | 60.12 | 88.00 | 46.77 | 52.03 | 95.00 |
Appendix F Additional Experimental Results and Analyses
F.1 Additional Experimental Results
@@
Main experiments. Table 8 presents the complete results of the main experiment. Each model is executed 5 times with varying random seeds, and the mean scores along with standard deviations are reported. “Avg. AUROC”, “Avg. FPR95”, “Avg. AUPRC”, and “Avg. Rank” indicate the average AUROC, FPR95, AUPRC, and rank across all datasets, respectively. We observe that the End-to-End approach stands out in AUROC, achieving the best overall results. However, in the metrics of AUPRC and FPR95, SSL-D in the two-step approach also performs well, showing competitive results.
@@
Robustness under Training Set Perturbation. While Observations ❼ and ❽ are based on AUROC, we also provide AUPRC and FPR95 results in this section. However, it is important to note that since AUPRC and FPR95 are computed based on the model’s performance on AUROC, the conclusions drawn from AUROC do not necessarily extend to these other metrics. As shown in Fig.9 and Fig.10, the results for AUPRC and FPR95 do not exhibit the same consistent trends, indicating that the behavior of models may vary across different evaluation metrics.
F.2 Additional Experimental Analyses
In this section, as illustrated in Fig. 2, we categorize the GLAD/GLOD methods into four main technical groups:
One-Class Classification: OCGIN, OCGTL.
Contrastive Learning: GLocalKD, GOOD-D, CVTGOD, SIGNET.
Graph Reconstruction: GLADC, GraphDE.
Well-Trained GNNs: AAGOD, GOODAT.
From Fig.11, we observe that contrastive learning methods generally perform better and exhibit more stable results compared to other methods. This performance difference can be attributed to several key factors:
Rich Representations in Contrastive Learning: Contrastive learning methods (GLocalKD, GOOD-D, CVTGOD, SIGNET) demonstrate superior and stable performance across all metrics, particularly AUROC and AUPRC. This strong performance can be attributed to their ability to maximize mutual information between different augmentations of the same graph (positive pairs) and minimize it for negative pairs. By capturing and preserving the most relevant and rich graph features, these methods produce robust, discriminative representations that generalize well in OOD detection. Their compact interquartile ranges indicate stable performance across datasets.
Sensitivity to Anomalies in One-Class Classification: One-class classification methods (OCGIN, OCGTL) perform well overall, particularly in FPR95. These methods excel by learning a dense, representative boundary around the in-distribution (ID) data, treating everything outside this boundary as OOD. Their strength lies in handling compact, well-defined distributions, making them particularly effective in scenarios where OOD samples deviate significantly from the ID distribution. This is why they show relatively consistent performance across datasets. However, when the OOD samples closely resemble the ID samples, the separation becomes more challenging, leading to some variability in performance on metrics like AUROC and AUPRC.
Limitations of Graph Reconstruction: Graph reconstruction methods (GLADC, GraphDE) show more variability. GLADC performs reasonably well on AUROC but struggles with AUPRC and FPR95. These methods reconstruct the input graph and detect anomalies based on reconstruction errors, but when the reconstruction error does not strongly correlate with OOD instances, their effectiveness decreases. This is especially true for GraphDE, which shows poor performance and significant variability.
Dependency on Pre-trained Models in Well-Trained GNNs: Well-trained GNNs (AAGOD, GOODAT) show mixed results, depending on the availability of pre-trained model parameters. These methods rely heavily on the pre-trained GNNs’ quality. GOODAT is more stable across datasets, while AAGOD displays higher variability. The methods perform well when high-quality pre-trained parameters are available, but retraining without these parameters leads to a noticeable performance drop, indicating a strong dependency on the initial pre-training phase.
Overall, contrastive learning methods stand out for their ability to capture high mutual information between graph augmentations, resulting in robust and generalizable representations. One-class classification methods perform well, especially in handling distinct OOD cases but show some sensitivity when OOD samples closely resemble ID data. Graph reconstruction methods face challenges in correlating reconstruction errors with OOD detection, leading to inconsistent results. Well-trained GNNs can be effective but are highly dependent on pre-trained models, impacting their consistency across different scenarios. This analysis highlights the importance of selecting methods that align with the characteristics of the task and data.
F.3 Judge Score Distribution Analysis
As illustrated in Fig.12, we visualize the OOD Judge score distributions for eight baseline models across three datasets (AIDS, BZR, Tox21_HSE, and AI-DH). The X-axis shows the OOD Judge score, and the Y-axis shows the frequency. These plots demonstrate how OOD and ID samples are distributed based on their scores, highlighting the separation between these distributions. This analysis directly assesses the models’ ability to distinguish OOD from ID samples, which is more task-relevant than visualizing the embedding space.
Most methods perform well on the AIDS dataset, with clear separation between OOD (red) and ID (blue) distributions. However, methods like GraphDE and SIGENT exhibit noticeable differences in distribution shapes, suggesting variability in decision boundaries and sensitivity to data characteristics.
In contrast, all methods show poor performance on the Tox21_HSE dataset, with overlapping OOD and ID distributions. This overlap indicates a failure to effectively distinguish OOD from ID samples, suggesting that the dataset presents significant challenges for OOD detection.
GK-D (two-step) | SSL-D (two-step) | GNN-based GLAD (end-to-end) | GNN-based GLOD (end-to-end) | |||||||||||||||
@@ F.3 Judge Score Distribution Analysis
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | AAGOD | GOODAT | |
p53 | ||||||||||||||||||
HSE | ||||||||||||||||||
MMP | ||||||||||||||||||
PPAR | ||||||||||||||||||
COLLAB | ||||||||||||||||||
IMDB-B | ||||||||||||||||||
REDDIT-B | ||||||||||||||||||
ENZYMES | ||||||||||||||||||
PROTEINS | ||||||||||||||||||
DD | ||||||||||||||||||
BZR | ||||||||||||||||||
AIDS | ||||||||||||||||||
COX2 | ||||||||||||||||||
NCI1 | ||||||||||||||||||
DHFR | ||||||||||||||||||
IM→IB | ||||||||||||||||||
EN→PR | ||||||||||||||||||
AI→DH | ||||||||||||||||||
BZ→CO | ||||||||||||||||||
ES→MU | ||||||||||||||||||
TO→SI | ||||||||||||||||||
BB→BA | ||||||||||||||||||
PT→MU | ||||||||||||||||||
FS→TC | ||||||||||||||||||
CL→LI | ||||||||||||||||||
HIV-Size | ||||||||||||||||||
ZINC-Size | ||||||||||||||||||
HIV-Scaffold | ||||||||||||||||||
ZINC-Scaffold | ||||||||||||||||||
IC50-Size | ||||||||||||||||||
EC50-Size | ||||||||||||||||||
IC50-Scaffold | ||||||||||||||||||
EC50-Scaffold | ||||||||||||||||||
IC50-Assay | ||||||||||||||||||
EC50-Assay | ||||||||||||||||||
Avg. AUROC | 52.69 | 53.67 | 57.56 | 52.91 | 52.11 | 50.35 | 61.29 | 61.50 | 66.00 | 64.29 | 73.14 | 75.86 | 65.50 | 71.07 | 71.05 | 60.38 | 61.54 | 61.91 |
Avg. Rank | 13.80 | 13.54 | 12.03 | 13.86 | 14.00 | 13.83 | 10.83 | 10.63 | 7.26 | 8.71 | 5.43 | 3.37 | 7.03 | 3.69 | 4.14 | 10.58 | 9.89 | 9.43 |
GK-D (two-step) | SSL-D (two-step) | GNN-based GLAD (end-to-end) | GNN-based GLOD (end-to-end) | |||||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | AAGOD | GOODAT | |
p53 | ||||||||||||||||||
HSE | ||||||||||||||||||
MMP | ||||||||||||||||||
PPAR | ||||||||||||||||||
COLLAB | ||||||||||||||||||
IMDB-B | ||||||||||||||||||
REDDIT-B | ||||||||||||||||||
ENZYMES | ||||||||||||||||||
PROTEINS | ||||||||||||||||||
DD | ||||||||||||||||||
BZR | ||||||||||||||||||
AIDS | ||||||||||||||||||
COX2 | ||||||||||||||||||
NCI1 | ||||||||||||||||||
DHFR | ||||||||||||||||||
IM→IB | ||||||||||||||||||
EN→MU | ||||||||||||||||||
AI→DH | ||||||||||||||||||
BZ→CO | ||||||||||||||||||
ES→MU | ||||||||||||||||||
TO→SI | ||||||||||||||||||
BB→BA | ||||||||||||||||||
PT→MU | ||||||||||||||||||
FS→TC | ||||||||||||||||||
CL→LI | ||||||||||||||||||
HIV-Size | ||||||||||||||||||
ZINC-Size | ||||||||||||||||||
HIV-Scaffold | ||||||||||||||||||
ZINC-Scaffold | ||||||||||||||||||
IC50-Size | ||||||||||||||||||
EC50-Size | ||||||||||||||||||
IC50-Scaffold | ||||||||||||||||||
EC50-Scaffold | ||||||||||||||||||
IC50-Assay | ||||||||||||||||||
EC50-Assay | ||||||||||||||||||
Avg. AUPRC | 43.27 | 40.20 | 46.52 | 45.51 | 47.85 | 48.53 | 55.11 | 54.87 | 62.15 | 56.96 | 64.98 | 66.24 | 56.21 | 61.77 | 62.54 | 49.97 | 57.89 | 56.71 |
Avg. Rank | 12.66 | 13.77 | 12.69 | 12.23 | 12.11 | 11.46 | 8.89 | 9.14 | 6.94 | 9.69 | 6.34 | 4.37 | 8.71 | 6.06 | 6.11 | 11.83 | 8.97 | 9.03 |
GK-D (two-step) | SSL-D (two-step) | GNN-based GLAD (end-to-end) | GNN-based GLOD (end-to-end) | |||||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | AAGOD | GOODAT | |
p53 | ||||||||||||||||||
HSE | ||||||||||||||||||
MMP | ||||||||||||||||||
PPAR | ||||||||||||||||||
COLLAB | ||||||||||||||||||
IMDB-B | ||||||||||||||||||
REDDIT-B | ||||||||||||||||||
ENZYMES | ||||||||||||||||||
PROTEINS | ||||||||||||||||||
DD | ||||||||||||||||||
BZR | ||||||||||||||||||
AIDS | ||||||||||||||||||
COX2 | ||||||||||||||||||
NCI1 | ||||||||||||||||||
DHFR | ||||||||||||||||||
IM→IB | ||||||||||||||||||
EN→PR | ||||||||||||||||||
AI→DH | ||||||||||||||||||
BZ→CO | ||||||||||||||||||
ES→MU | ||||||||||||||||||
TO→SI | ||||||||||||||||||
BB→BA | ||||||||||||||||||
PT→MU | ||||||||||||||||||
FS→TC | ||||||||||||||||||
CL→LI | ||||||||||||||||||
HIV-Size | ||||||||||||||||||
ZINC-Size | ||||||||||||||||||
HIV-Scaffold | ||||||||||||||||||
ZINC-Scaffold | ||||||||||||||||||
IC50-Size | ||||||||||||||||||
EC50-Size | ||||||||||||||||||
IC50-Scaffold | ||||||||||||||||||
EC50-Scaffold | ||||||||||||||||||
IC50-Assay | ||||||||||||||||||
EC50-Assay | ||||||||||||||||||
Avg. FPR95 | 94.02 | 92.76 | 88.88 | 93.54 | 89.39 | 89.02 | 77.13 | 76.75 | 75.66 | 76.60 | 66.40 | 69.62 | 74.59 | 73.72 | 75.02 | 91.23 | 83.45 | 81.59 |
Ave. Rank | 13.17 | 12.03 | 13.09 | 12.69 | 11.77 | 11.63 | 8.43 | 7.46 | 8.17 | 9.66 | 6.57 | 4.54 | 7.29 | 7.09 | 7.20 | 12.03 | 9.54 | 8.66 |